Find the foot of the perpendicular drawn from the point A(1,0,3) to the join of the points B(4,7,1) and C(3,5,3)

Grade:Upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
Let ‘L’ be the fot of perpendicular AL, which divides BC in the ratio a:1 =
(3a+4)/(a+1),(5a+7)/(a+1),(3a+1)/(a+1)
Also D.R’s of the BC and AL are ,
D.R’s of BC= -1,-2,2
D.R’s of AL =((3a+4)/(a+1))-1,((5a+7)/(a+1))-0,((3a+1)/(a+1))-3
since AL is perpendicular to BC
Cos90 = 0 →BC, AL are the lines with D.R’s as magnitudes
2(((3a+4)/(a+1))-1)-2(((5a+7)/(a+1))-0)+(((3a+1)/(a+1))-3)2=0
we get a=-7/4
So L=(5/3,7/3,17/3)